OpenStudy (anonymous):

When solving an equation for x, if 5 is added to one side of the equation, what needs to be done to the other side of the equation? A. add 5 B. divide by 5 C. subtract 5 D. nothing

OpenStudy (shamil98):

you add 5 to the other side as well

OpenStudy (shamil98):

x - 5= 0 +5 +5 x= 5 for example.

OpenStudy (anonymous):

yeah shamil98 is right

OpenStudy (anonymous):

What you do to one side you have to do for the other

OpenStudy (anonymous):

thanks :)